OpenStudy (anonymous):

-4x^-3y^2z^-4/2x^2y^-1z^-6 Please help.....I'm getting 2 solutions and not sure which is correct. b) -2y^3z^2/x^5 c) y^2/-2x^5z^10

OpenStudy (anonymous):

That's incorrect! But thanks for contributing :). However, I'm confident the solution is one of my solutions (b) or (c), just not sure which.

OpenStudy (anonymous):

so much powers

OpenStudy (anonymous):

the question is not so clear......

OpenStudy (anonymous):

The equation is: -4x^-3y^2z^-4/2x^2y^-1z^-6 -4x^-3y^2z^-4 (divided by) 2x^2y^-1z^-6

OpenStudy (campbell_st):

break the problem up its \[\frac{-4}{2} x^{-3-2}y^{2 - -1} z^{-4 - -6}\] which simplifes to \[-2x^{-5} y^3x^2 = \frac{-2y^3z^2}{x^5}\]

OpenStudy (anonymous):

Thank you!

OpenStudy (phi):

These are not that hard if you remember that x^2 means x*x and x^3 means x*x*x so if you have a question like: \[ \frac{x^2}{x^1}= \frac{x \cdot x}{x}= \frac{\cancel{x}\cdot x}{\cancel{x}} =x^1 \] to remind yourself it is the top exponent minus the lower exponent.